what is the height os triangl of base of 16inches and area of 32in square
Answer by ewatrrr(24785)


Hi
A(triangle) = (1/2)b*h
Question states***
32in^2 = (1/2) 16*h
64/16 = h
4in = h
CHECKING our Answer***
(1/2)b*h =(1/2)16in*4in = 32in^2